# Uni-NaVid: A Video-based Vision-Language-Action Model for Unifying Embodied Navigation Tasks

### v1

@@ 3.1 Model of Uni-NaVid
3.2 Training of Uni-NaVid
@@
To train a large-scale model for generalist navigation, it is essential to gather extensive and diverse navigation data across various tasks and environments. However, directly collecting vast amounts of real-world navigation data can be cost-prohibitive. To address this, we propose two key approaches for training Uni-NaVid: first, we collect multi-task navigation data in a wide range of synthetic environments (861 scenes in total), enabling Uni-NaVid to acquire general navigation skills; second, we co-tune Uni-NaVid with real-world video-based question-answering data, enhancing its understanding of real-world images and supporting its open-vocabulary knowledge.
@@
Multi-task navigation data. We collect the largest multi-task navigation dataset to date on the Habitat simulator environment [64], including 3.6M samples across four different navigation tasks introduced below. They are all curated in a unified format. Details are in Supplementary.
(A) Vision-and-language navigation [37] requires the agent to comprehend language instructions in order to follow the described path. We collect two types of instructions: one type uses landmarks and describes required motions between them, while the other specifies low-level actions, such as moving forward by 5 steps. We collect 2.4M samples based on both VLN-CE R2R [37] and RxR [40].
@@
(B) Object goal navigation [64]  involves the agent navigating an environment to locate a specific object based on provided visual or linguistic cues. This task assesses the agent’s ability to perceive objects, understand spatial relationships, and execute efficient search strategies. We gather 483k samples from datasets on Habitat Matterport 3D dataset [57].
@@
(C) Embodied question answering [18] requires the agent to navigate to the related area for question answering. It involves spatial reasoning, object description, and understanding contextual information, requiring the ability to integrate perception, language comprehension, and decision-making. Following the setup in EQA [18], the agent first navigates to the target related to the question, issues a stop action, and then provides an answer. We collect 240k video-action samples and 10k video-answering samples on the EQA dataset [18] on Matterport 3D environments [10].
@@
(D) Human following [31] requires the agent to track and follow a human target with a specific description in dynamic and crowded environments, e.g., “Follow the man in the blue t-shirt.”. The agent must recognize the appearance of the human, follow the correct person described in the instructions, predict their movement trajectory, and keep an appropriate distance while avoiding obstacles [20]. Based on the Habitat 3.0 social navigation environment [56], we collect 544k human-following navigation samples.
The data statistics can be found in Figure 3. Note that all data were collected from synthetic environments, including Habitat-Matterport and Matterport 3D, totaling 861 scenes. We use the default settings of each environment, with a height range from $0.88$ m to $1.25$ m and a robot radius between $0.1$ m and $0.6$ m. This helps avoid overfitting to a specific robot embodiment.
Joint training on synthetic and real-world data. Although we collect navigation data from various environments, the diversity in both observations and instructions remains limited to a certain set of synthetic environments. To incorporate open-world knowledge, we follow previous VLA methods [95, 8], fine-tune Uni-NaVid using open-world video question-answering datasets. We collect 2.3M video question-answering data from publicly available datasets [7, 16, 44], leading to a total of 5.9M training samples. We adopt a two-stage training process: train the visual projector first, and then jointly train both the projector and the large language model. Further training details are provided in the Supplementary Materials.
3.3 Implementation Details
@@
Training configuration. Uni-NaVid is trained on a cluster server with 40 NVIDIA H800 GPUs for approximately 35 hours, totaling 1400 GPU hours. For video data, we sample frames at 1 FPS to remove redundant information between consecutive frames. During training, EVA-CLIP [70] and Vicuna-7B [17] are pre-loaded with default pre-trained weight. Following the training strategy of VLM [46], we optimize the trainable parameters for only 1 epoch.
Benchmark evaluation and real-world deployment. For each navigation task, we adhere to its default settings [37, 64, 18, 31]. Specifically, for EQA [18], the agent executes actions based on the policy until a stop command is output. We then remove the navigation special token <NAV> and query the questions using the navigation history. For real-world deployment, we use a remote server with an NVIDIA A100 GPU to run Uni-NaVid, which processes observations (along with text instructions) and sends commands to a local robot to execute the predicted actions. Uni-NaVid requires approximately 0.2 seconds to generate the next four actions. During navigation, the robot asynchronously uploads the latest observations to the model while executing and pending future actions. Refer to the supplementary video for real-world navigation performance.
4 Experiment
@@
We conduct experiments to evaluate Uni-NaVid on three specific aspects: (1) How does Uni-NaVid perform on individual tasks? (2) Does learning multiple navigation tasks lead to synergistic improvements? (3) Is the key design of our method effective? Due to space limitations, the complete set of protocols, additional results, and visualizations are provided in the Supplementary Materials.
Benchmarks. We evaluate our method on various benchmarks across different navigation tasks. For vision-and-language navigation tasks, we use Val-Unseen splits of R2R [37] and RxR [39] benchmarks. For object goal navigation, we use the validation split of Habitat Matterport 3D dataset [57]. For the embodied question-answering task, we use validation split of MP3D-EQA benchmark [18]. For the human following task, we leverage the Habitat 3.0 social navigation benchmark [56] and extend it to multi-person environments. For video understanding, we follow the evaluation manner of existing methods [44]. We choose ScanQA [7], MSVD [12], MSRVTT [81], and ActivityNet [9] for video question-answering evaluation.
@@
4.1 Individual Task Results
@@
Table 3: Vision-and-language navigation (RxR). Comparison on VLN-CE RxR [40] Val-Unseen. : only trained on VLN-CE R2R.
Comparison on vision-and-language navigation. We evaluate our method with mainstream baselines on two publicly available benchmarks: VLN-CE R2R [37] and RxR [40]. The results are shown in Table 2 and Table 3. We find that our methods achieve SOTA-level performance on both datasets using only RGB videos as observations. In comparison to NaVid [95], which is also a vision language model that is solely trained on VLN data, our approach demonstrates significant improvements, with a $+25.7\%$ increase in Success Rate (SR) on R2R. For zero-shot methods (InstructNav [51] and A2Nav [15]) that use ChatGPT with only text inputs for visual language navigation (VLN), these approaches often face challenges in transitioning between text prompts and visual information, resulting in less than satisfactory outcomes (almost half the SR of Uni-NaVid).
@@
Table 4: Object goal navigation. Comparison on Habitat Matterport 3D [57] ObjectNav dataset.
Comparison on object goal navigation. We conduct the experiments on HM3D [57] to compare Uni-NaVid with mainstream methods [77, 58, 59, 84, 83] that also learn from ObjectNav data. The results, shown in Table 4, demonstrate that our approach achieves the best performance. Note that methods not utilizing odometry face challenges as they must rely on implicit memory to retain the historical trajectory. Nevertheless, Uni-NaVid still achieves significant gains in SR (+$4.7\%$) and SPL (+$8.8\%$) compared to previous state-of-the-art methods. Additionally, we believe our method’s ObjectNav performance can be further enhanced by incorporating reinforcement learning techniques, as demonstrated by PIRLNav [59] and Poliformer [90].
@@
Table 5: Embodied question answering. Comparison on Habitat Matterport3D EQA dataset [18].
Comparison on embodied question answering. The evaluation results on MP3D-EQA [18] are presented in Table 5. Despite navigating in continuous environments (CE), our method outperforms existing approaches (e.g., NaviLLM [98] leverage the same evaluation strategy in Sec. 3.3) that operate within discrete landmark-based environments (DE). Moreover, when provided with the ground truth (GT) navigation trajectory, our method shows a significant improvement, demonstrating its ability to understand navigation history effectively.
@@
Comparison on human following. We compared our method with two most relative methods PoliFormer [90] and IBVS [23]. Since both methods require a specific human bounding box as input, obtained from an upstream algorithm, we use the bounding box from the open-world object detector GroundingDINO [49] and the ground truth provided by the simulator to evaluate the human following performance of the comparison methods under various setups. As shown in Table 6, Uni-NaVid outperforms the comparison methods on both SR (+$21.0\%$) and FR (+$4.4\%$) while maintaining low CR under any setup, even when they use ground truth bounding boxes as input. This demonstrates that Uni-NaVid can effectively infer instructions and follow the correct human, as well as predict the human’s movement patterns accurately.
Figure 4: Visual results of Uni-NaVid in real-world. We deploy Uni-NaVid across diverse environments to execute instructions in a zero-shot setting. We provide both third-person views and the robot’s trajectory, showing effective multi-task navigation performance.
@@
Table 7: Embodied video question answering. Comparison on ScanQA [7] benchmark.
@@
Table 8: Video question answering. Comparison with leading methods (all based on Vicuna-7B [17]) on VQA benchamarks.
@@
Comparison on video question answering. We first evaluate our method on ScanQA [7] on Tab. 7. Compared to mainstream baselines, we find that Uni-NaVid archives the best performance on four metrics, including BLUE-1 ($+17.9\%$), ROUGE ($+5.7\%$), METEOR ($+16.2\%$), and CIDEr ($+13.1\%$). This proves the superiority of our methods on spatial scene understanding. Note that the EM metric requires an exact match between the question and answer, which is not well-suited to our method, as it is designed to learn from diverse data and generate flexible responses. We also compare our method on open-ended video question answering benchmark in Tab. 8, where we show comparable performance to SOTA methods.
4.2 Qualitative Results in Real-World
Figure 4 visualizes the results of all tasks in real-world environments. Despite training only on synthetic data, Uni-NaVid demonstrates impressive adaptability across various real-world settings, successfully performing multi-task navigation. It effectively aligns navigation history with given instructions, identifies novel objects outside the training set, and avoids obstacles in densely populated areas. Additionally, it reliably follows human movement across diverse environments. Supplementary materials provide numerical results and further visualizations of challenging scenarios.
Figure 5: Comparsion on data scale and multi-task training.
4.3 Ablation Study

### v2

@@ IV-C Action Planning
Input: $\{Long\_term\_tokens\}\{Shot\_term\_tokens\}\\ \{Current\_tokens\}<\!NAV\!>\{Instruction\}$ Output: $<\!Action\_0\!>\!<\!Action\_1\!>\!<\!Action\_2\!>\!\\ \!<\!Action\_3\!>$
[…]
V Data Collection and Training
To train  Uni-NaVidfor mastering multi-navigation tasks, it is crucial to gather extensive and diverse navigation data across various tasks and environments. However, directly collecting large amounts of real-world navigation data can be prohibitively expensive. To address this challenge, we propose two key strategies for training Uni-NaVid: First, we collect multi-task navigation data from a wide range of synthetic environments (totaling 861 scenes) using a uniform input and output format, enabling Uni-NaVidto acquire general navigation skills. Second, we co-tune Uni-NaVidwith real-world video-based question-answering data, enhancing its ability to interpret real-world images and supporting its open-vocabulary knowledge acquisition.
@@ V Data Collection and Training
V-A Multi-Task Navigation Data.
@@
We collect the largest multi-task navigation dataset to date within the Habitat simulator environment [76], comprising 3.6 million samples across four distinct navigation tasks, as described below. All tasks are curated within a unified framework. A detailed data collection strategy is provided in the Supplementary Materials.
(A) Vision-and-language navigation [44, 46] require the agent to interpret and ground instructions in visual observations, effectively combining linguistic and visual information to make sequential decisions. Specifically, the agent has to navigate based on landmarks and motions described in the text and stop nearby the correct destination. Here, we collect 2.4M navigation samples of mainstream VLN datasets, VLN-CE R2R [44] and RxR [47], that focus on continuous environments.
@@
(B) Object Goal Navigation [76] involves an agent navigating an environment to locate a specific object based on provided visual or linguistic cues. This task evaluates the agent’s ability to perceive objects, understand scene layout, and execute efficient search strategies. We collected 483k samples from datasets in the Habitat Matterport 3D dataset (HM3D ObjectNav) [70]. Note that, in HM3D ObjectNav, the agent is required to locate objects from a predefined category set (e.g., sofa, chair, and bed). Nevertheless, experiments demonstrate that our method generalizes to SOTA-level open-vocabulary object goal searching, as shown in Table V.
@@
(C) Embodied question answering [90] requires the agent to navigate to the related area for question answering. It involves spatial reasoning, object description, and understanding contextual information, requiring the ability to integrate perception, language comprehension, and decision-making. Following the setup in main stream EQA methods [21, 90], the agent first navigates to the target related to the question, issues a stop action, and then provides an answer. We collect 240k video-action samples and 10k video-answering samples on the MP3D-EQA dataset [21] on Matterport 3D environments [11].
@@
(D) Human following [37, 25] requires the agent to track and follow a human target with a specific description in dynamic and crowded environments, e.g., “Follow the man in the blue t-shirt.”. The agent must recognize the appearance of the human, follow the correct person described in the instructions, predict their movement trajectory, and keep an appropriate distance while avoiding obstacles.
Figure 4: Language-described human following benchmark. We construct our human-following benchmark based on Habitat 3.0 [68] by incorporating textual descriptions for each avatar (eight in total, top row). The robot is required to comprehend these descriptions and accurately follow the designated individual in crowded environments.
However, there is currently no human-following dataset that supports language-described human following in crowded environments (multi-person scenarios). To this end, we extend the Habitat 3.0 social navigation benchmark [68] by (1) adding textual descriptions for each avatar (8 in total, as illustrated in Fig. 4), (2) introducing additional distracting human avatars to simulate challenging real-world environments, and (3) deploying the robot and humans in the Habitat Matterport 3D dataset [101], which offers photo-realistic rendering quality and diverse large-scale scenes. The robot and target human are initialized nearby (using the same setting as  [68]), with randomly moving distracting human avatars. Based on this setup, we collected 544k human-following navigation samples. We also add a detailed description in Supplementary Material. This benchmark will also be released to benefit the navigation community.
Unified navigation samples. The data statistics are presented in Figure 3. It is worth noting that the number of samples in VLN is relatively larger compared to other tasks. This is because VLN [44, 46] requires the agent to navigate all landmarks described in the instructions, which often results in longer trajectories and, consequently more video-action samples. Here, we collect all navigation samples in a uniform format, including an egocentric RGB video, a natural language instruction, and four corresponding future actions. All data were collected from synthetic scenes across the Habitat-Matterport 3D (HM3D) and Matterport 3D (MP3D) datasets. We use the default settings of each environment, with a height range of $0.88$ m to $1.25$ m and a robot radius between $0.1$ m and $0.6$ m. This approach helps prevent overfitting to a specific robot embodiment. This approach helps prevent overfitting to a specific robot embodiment. Note that while there exist insightful techniques [24, 29] investigating navigation for robots of general sizes, our focus is primarily on uniform multi-task navigation.
V-B Training Strategy of Uni-NaVid
[…]
@@ V-B Training Strategy of Uni-NaVid
Training configuration. Uni-NaVid is trained on a cluster server with 40 NVIDIA H800 GPUs for approximately 35 hours, totaling 1400 GPU hours. For video data, we sample frames at 1 FPS to remove redundant information between consecutive frames. During training, the vision encoder (EVA-CLIP [82]) and large language model (Vicuna-7B [20]) are pre-loaded with default pre-trained weight. Following the training strategy of VLM [54], we optimize the trainable parameters for only 1 epoch.
VI Experiment
@@
We conduct experiments to evaluate Uni-NaVid on three specific aspects: (1) How does Uni-NaVid perform on individual tasks? (2) Does learning multiple navigation tasks lead to synergistic improvements? (3) Is the key design of our method effective? To evaluate the general-purpose navigation method, we conduct extensive experiments on individual navigation tasks, employing corresponding strong baselines. Additional details are provided in the supplemental material.
Benchmarks. We evaluate our method on various benchmarks across different navigation tasks. Given the diversity of benchmarks spanning various environments and simulators, we meticulously verify the scene splits to ensure no overlap exists between the training and validation scenes across benchmarks.
  • Vision-and-language navigation: We test our method on the validation splits of the VLN-CE R2R [44] and RxR [46] benchmarks.
  • Object goal navigation: We use the validation split of the Habitat Matterport 3D (HM3D) dataset [70], which requires the agent to find target objects from six categories (sofa, chair, TV, bed, toilet, and plant) in unseen environments. Moreover, to test generalizability, we also evaluate our method on the HM3D-OVON dataset [101], an open-vocabulary object navigation benchmark, in a zero-shot manner.
  • Embodied question-answering: We use the validation split of the MP3D-EQA benchmark [90]. Additionally, we conduct experiments on the more recent Embodied Video Question Answering benchmark, OpenEQA [63].
  • Human following: We evaluate our method alongside mainstream approaches on our proposed language-described human following benchmark.
  • Video understanding: We follow the evaluation procedures of existing VQA methods [51]. We choose the ScanQA [7], MSVD [13], MSRVTT [96], and ActivityNet [10] datasets.
@@
VI-A Deployment Details of Uni-Navid.
Benchmark evaluation. For each navigation task, we adhere to the default settings of each navigation task [44, 76, 21, 37]. All tasks take an online captured RGB video (capturing one frame after each action) and a textual instruction as inputs, and output the next four actions (Sec. IV-C). The robot then executes the predicted actions and calls STOP once the first predicted action is a stop action. For VLN and EQA tasks, we directly use the text instruction provided by the benchmark episodes. For human following and object goal navigation, we transform the target information into an instruction by adding prefixes such as ”Search for” or ”Follow.” Further details can be found in the supplemental material.
It is worth noting that for EQA [21] task, the agent executes navigation actions until a stop command is issued. We then remove the navigation-specific token <NAV> and query the questions using the navigation history. This strategy alleviates the ambiguity for the LLM in deciding whether to navigate or answer a question (See Table X).
Real-world deployment. For real-world deployment, we utilize a remote server with an NVIDIA A100 GPU to run Uni-NaVid, which processes observations (along with text instructions) and sends commands to a local robot to execute the predicted actions. Uni-NaVid requires approximately 0.2 seconds to generate the next four actions. During navigation, the robot asynchronously compresses and uploads the latest observations to the model while executing pending actions. Refer to the supplementary video for real-world navigation performance.
VI-B Individual Task Results
@@
TABLE III: Vision-and-language navigation (RxR). Comparison on VLN-CE RxR [47] Val-Unseen. : only trained on VLN-CE R2R.
[…]
@@
TABLE IV: Object goal navigation. Comparison on Habitat Matterport 3D [70] ObjectNav dataset.
Comparison on object goal navigation. We conduct the experiments on HM3D [70] to compare Uni-NaVid with mainstream methods [91, 71, 72, 99, 98] that also learn from ObjectNav data. The results, shown in Table IV, demonstrate that our approach achieves the best performance. Note that methods not utilizing odometry face challenges as they must rely on implicit memory to retain the historical trajectory. Nevertheless, Uni-NaVid still achieves significant gains in SR (+$4.7\%$) and SPL (+$8.8\%$) compared to previous state-of-the-art methods. Additionally, we believe our method’s ObjectNav performance can be further enhanced by incorporating reinforcement learning techniques, as demonstrated by PIRLNav [72] and Poliformer [106].
Method VAL SEEN VAL SEEN SYNONYMS VAL UNSEEN
[…]
BC 11.1 4.5 9.9 3.8 5.4 1.9
DAgger 11.1 4.5 9.9 3.8 5.4 1.9
RL 18.1 9.4 15.0 7.4 10.2 4.7
BCRL 39.2 18.7 27.8 11.7 18.6 7.5
DAgRL 41.3 21.2 29.4 14.4 18.3 7.9
VLFM [100] 35.2 18.6 32.4 17.3 35.2 19.6
DAgRL+OD [101] 38.5 21.1 39.0 21.4 37.1 19.8
Uni-NaVid 41.3 21.1 43.9 21.8 39.5 19.8
TABLE V: Object goal navigation. Comparison on HM3D-OVON [101]. ${}^{*}:$ denotes zero-shot methods.
To evaluate the generalization ability for open-vocabulary objects, we evaluate our method on the open-vocabulary object goal navigation benchmark (HM3D-OVON [101]) in a zero-shot manner. The results in Table V demonstrate that our method achieves significant improvement over the zero-shot method (VLFM [100]) and even outperforms the fine-tuned method (DAgRL+OD [101]) on the VAL SEEN and VAL UNSEEN splits. This proves the generalizability of our method.
@@
TABLE VI: Embodied question answering. Comparison on Habitat Matterport3D EQA dataset [21].
Comparison on embodied question answering. The evaluation results on MP3D-EQA [90] are presented in Table VI. Despite navigating in continuous environments (CE), our method outperforms existing approaches (e.g., NaviLLM [114] leverage the same evaluation strategy in Sec. VI-A) that operate within discrete landmark-based environments (DE). Moreover, when provided with the ground truth (GT) navigation trajectory, our method shows a significant improvement, demonstrating its ability to understand navigation history effectively. We also report our performance on the more challenging EM-EQA benchmark, OpenEQA [63], in the Supplemental Material, which includes more complex questions. Our method achieves comparable performance to GPT-4V with scene captions [63].
@@
Comparison on human following. We compared our method with two most relative methods PoliFormer [106] and IBVS [28]. Since both methods require a specific human bounding box as input, obtained from an upstream algorithm, we use the bounding box from the open-world object detector GroundingDINO [58] and the ground truth provided by the simulator to evaluate the human following performance of the comparison methods under various setups. As shown in Table VII, Uni-NaVid outperforms the comparison methods on both SR (+$21.0\%$) and FR (+$4.4\%$) while maintaining low CR under any setup, even when they use ground truth bounding boxes as input. This demonstrates that Uni-NaVid can effectively infer instructions and follow the correct human, as well as predict the human’s movement patterns accurately. We include additional human-following experiments in various environments, such as HSSD [42] and MP3D [11], in the Supplemental Material. Our method consistently demonstrates SOTA performance across these settings.
Figure 5: Visual results of Uni-NaVid in real-world. We deploy Uni-NaVid across diverse environments to execute instructions in a zero-shot setting. We provide third-person views with robot’s trajectory, showing effective navigation performance. We indicate the starting point as a blue dot and the ending point as a green arrow.
@@
TABLE VIII: Embodied video question answering. Comparison on ScanQA [7] benchmark.
@@
TABLE IX: Video question answering. Comparison with leading methods (all based on Vicuna-7B [20]) on VQA benchamarks.
Figure 6: Vivusal results of Uni-NaVid on compostional tasks. The agent is required to execute complex instructions involving multiple navigation tasks. Our method successfully accomplishes these navigation tasks sequentially. Notably, both the instructions and environments are novel to our approach. Please refer to the supplementary videos.
@@
Comparison on video question answering. We first evaluate our method on ScanQA [7] on Tab. VIII. Compared to mainstream baselines, we find that Uni-NaVid archives the best performance on four metrics, including BLUE-1 ($+17.9\%$), ROUGE ($+5.7\%$), METEOR ($+16.2\%$), and CIDEr ($+13.1\%$). This proves the superiority of our methods on spatial scene understanding. Note that the EM metric requires an exact match between the question and answer, which is not well-suited to our method, as it is designed to learn from diverse data and generate flexible responses.
We further evaluate our method on open-ended video question-answering benchmarks [13, 96, 10], as presented in Table IX. To ensure a fair comparison, we focus on methods that employ the same large language model backbone (Vicuna-7B [20]). The results indicate that even after extensive token merging (Sec. IV-B), Uni-NaVidachieves performance comparable to state-of-the-art methods. This demonstrates the effectiveness of both our token merging and training strategies, while also highlighting robust open-world understanding capabilities.
VI-C Qualitative Results in Real-World
We conducted extensive experiments on real-world environments (experiment details are provided in the supplemental material) under diverse environments in a zero-shot manner. Notably, both the instructions and environments are novel to our method. We first evaluated the performance of individual navigation tasks (Fig. 5), including (A) vision-and-language navigation, (B) object goal navigation, (C) embodied question answering, and (D) human following. We found that Uni-NaVid can understand diverse instructions and demonstrates impressive performance in long-horizon navigation tasks (e.g., navigating across hallways and entering rooms), as well as in searching for out-of-vision objects and answering subsequent questions. Moreover, the agent is capable of following a human even when the person’s appearance deviates from the description of the avatar in the human-following dataset (Sec. V-A). The statistics of the corresponding real-world experiments can be found in the Supplemental Material.
In addition to individual navigation tasks, we also evaluate our method on more complex instructions involving multiple navigation tasks (Fig. 6). In this scenario, the agent is required to sequentially execute the navigation tasks described in the language instructions. Our model demonstrates impressiove performance in aligning the current navigation process with the instructions to reason about the current state of navigation. Furthermore, we provide a detailed illustration of action prediction during navigation in Fig. 7, where we plot the predicted action probabilities of Uni-NaVid. Notably, with only slight differences in object descriptions, e.g., ’chair with a toy’ and ’chair with a sweater’. Specifically, our method successfully distinguishes between the locations and predicts actions accordingly. Interestingly, the action probabilities (for the next four actions) reveal a sequential order of actions: first turning right/left, followed by moving forward. We provide additional visual results of our method in the Supplemental Material and encourage the audience to view our video, which showcases the real-world performance of our method.
Figure 7: Action prediction on the VLN tasks. We evaluate Uni-NaVid on challenging open-vocabulary objects, requiring it to recognize the target objects and follow the specified motions. We provide the predicted action probabilities (for the next four actions) to demonstrate its break-in navigation capability.
Figure 8: Comparsion on multi-task training and data scale. (a) We present the multi-task synergy of our method, illustrating the performance comparison between training with a single task and training with multiple tasks; (b) we demonstrate the performance across different navigation tasks under varying numbers of training samples.
VI-D Ablation Study